Need a simple solution and please don't copy from other solutions.
2.3-11. If the moment-generating function of X is 2 find the mean, variance, and pmf of X
Answer #1

The moment generating function of the random variable X is defined as follows:

M(t) = E(etx )

For discrete random variable, we have

M(t) = E(etx ) = P(X = 0)e0t + P(X = 1)e1t + P(X = 2) e2t + P(X = 3)e3t+...

Comparing this mgf with the given , then we get the following probability distribution of X.

X p(x)
1 2/5
2 1/5
3 2/5
Total 1

This is the pmf of X.

Let's find mean and variance of X.

Mean = E(X)

E(X)=sum Xp(x)=1*(2/5)+2*(1/5)+3*(2/5)=10/5=2

Therefore mean of X is 2

Variance of X = V(X)

V(X) = 〉 p(z)(X-E(X))2 (2/5)(1-2)2+11/5)(2-2)2+(2/5)(3-2)

(2/5) + 0 + (2/5)-4/5-0.80

Therefore, variance of X is 0.80